# Observation of an Accreting Planetary-Mass Companion with Signs of Disk–Disk Interaction in Orion ††thanks: Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programme ID 114.28JY

### v1

@@ Appendix A Data presentation and reduction
We use the JWST NIRCam images of the ONC from the PDRs4All Early Release Science (ERS) program (Berné et al. 2022, PID 1288) that were previously presented in Habart et al. (2024). We select the short-wavelength channel (i.e., for m), which includes seven filters (F140M, F162N, F164N, F182M, F187N, F210M, and F212N) with a 0.03” angular resolution. Details on the data reduction procedure to obtain these images are given in Habart et al. (2024). The images themselves are available through the specific PDRs4All MAST archive111https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html and through the REGARDS database222https://regards.osups.universite-paris-saclay.fr/user/jwst. We also use the Hubble Space Telescope (HST) images from the ACS/WFC of V2376 Ori presented in Ricci et al. (2008). This includes five filters (F435W, F555W, F658N, F775W, and F850LP), covering a wavelength range from 0.435 to 0.85 m.
The VLT MUSE observations were taken on the Feb 27, 2025 (DDT P114 proposal 114.28JY). The observations originally consisted of four dithers, each with an integration time of 1800 sec and 90 degrees rotation between them. Due to a technical loss only the first dither position was successfully observed, yet, after inspection, the quality of this single exposure was deemed sufficient for this initial analysis.
The data is reduced using the ESO MUSE data reduction pipeline (v.2.10.14) together with the python wrapper, which is part of MUSEpack (Zeidler 2019; Zeidler et al. 2019). We apply all steps of the the MUSE data reduction pipeline using the ESO provided master calibrations (MASTERBIAS, MASTERDARK, MASTERFLAT). To subtract the background we use the standard method “model” using 10% of the datacube spaxels (fraction=10) after ignoring the darkest 10% (ignore=10). We visually inspect the skymask to ensure that it selects the darkest regions of our cube while rejecting all artifacts. Furthermore, we carefully inspect the sky spectrum and sky continuum to ensure that only telluric lines are removed from the final datacube. Once all data will be available we will combine these with the current dataset, which will further enhance our data product by allowing us to remove any remaining artifacts and significantly increase the signal-to-noise.
@@ B.1 Relative position between HST and JWST images
We analyze the relative position of V2376 Ori and the faint source at two different epochs using the HST F435W image (taken in October 2004) and JWST F187N (taken in September 2022) images. We exclude the MUSE image from this analysis because there are no other sources in the MUSE field of view, which prevents astrometric corrections. Fig. 4 shows the contours of the HST F435W (blue continuous lines) and JWST F187N (red dashed lines) images. Both are centered on V2376 Ori, and oversampled on a common grid to ensure alignment to about a tenth of a NIRCam pixel (0.003”). The faint source (to the southwest of V2376 Ori) seems to have moved = 9 mas, dec = 25 mas between the two epochs, i.e., over nearly 18 years. Considering that the pixel size is 50 mas for Hubble and 30 mas for NIRCam, this difference is compatible with the absence of relative movement of the two sources over this time period.
@@ B.2 Angular separation of the sources
We estimate the angular separation between V2376 Ori and its companion in the HST F435W image, JWST F164N image and in the integrated MUSE cube. We first extract an intensity profile along a cut performed along the axis connecting the two sources. We then determine the position of V2376 Ori and its companion in the profile with the highest and second highest intensity peak, respectively. The distance in pixels between these two peaks is finally converted into an angular separation considering the respective pixel sizes in arcsec of the different instruments333the pixel sizes for ACS/WFC HST, NIRCam JWST and VLT MUSE data are respectively 0.05”, 0.031”, and 0.025”. This yields for the F435W HST image, F164N JWST image and the MUSE image an angular separation of 0.230.05, 0.250.03” and 0.230.03”, respectively. The final angular separation is considered as the mean of these values, leading to = 0.230.03”.
@@ Appendix C Extraction of the Spectral Energy Distribution
Fig.3 shows the spectral energy distribution (SED) of V2376 Ori b extracted from HST-ACS/WFC and JWST-NIRCam images. The flux arising from V2376 Ori b is contaminated by the background nebular emission from the ONC (see Fig. 1 C) and by the diffraction spikes of V2376 Ori (see, e.g., Fig. 1 F, G). The nebular emission is estimated using the Background2D function from the Photutils Python library444https://photutils.readthedocs.io/en/2.1.0/user_guide/index.htm. We use 5050 pixel boxes to create a low-resolution map with a sigma-clipping of three, which is then filtered with a 2D median filter of 55 pixels and interpolated to obtain the final 2D background. The images are then subtracted from this 2D background. Next we use three different approaches to remove the contribution of the diffraction spikes:
@@
For each image, we subtract a version rotated by 180° around the stellar center, so that every diffraction spike is canceled by its counterpart on the opposite side of the star.
@@
Finally, the on-source flux in each background-subtracted and PSF corrected image is extracted using the AperturePhotometry function from the Photutils library with an aperture radius of 0.08” centered on V2376 Ori b, which is shown in Fig. 1 (G) with the red dashed circle. The final flux in a given filter is the mean of the fluxes obtained for the different PSF subtraction methods.
@@ Appendix D Physical properties of the V2376 Ori system
Table 1 summarizes the physical properties of the companion V2376 Ori b. Fig. 5 shows the luminosity (top panel) and temperature (bottom panel) of V2376 Ori b, used to estimate its mass according to evolutionary tracks from Baraffe et al. (2003). The derivation of the companion’s luminosity is described in Sect. 3.1.
@@ E.1 Extraction of MUSE spectra
The spectrum of V2376 Ori and its companion is extracted from the MUSE cube using, respectively, a 4-pixels and a 3.15-pixels aperture radius, centered on each source coordinates. The smaller aperture for the companion is chosen to limit contamination from the host star. The background emission, which is dominated by the nebular emission from the ONC, is estimated following the method used by Eriksson et al. (2020) for Delorme 1 (AB)b, with two apertures of the same radius as the source, placed on either side of the source, in the direction perpendicular to the axis star-companion. The mean of the two obtained background spectra is then subtracted from the source spectrum. The direct subtraction leads to artifacts with negative intensities and P-cygni-like profiles, possibly produced by velocity effects and/or thermal broadening. We apply a gaussian filter with a standard deviation of 2.5 pixels to remove these effects. The obtained spectra are then calibrated using the photometric fluxes derived in the same aperture in the F775W HST filter using the MPDAF Python package and associated filter throughput. This filter’s wavelength range is fully covered by MUSE and is not dominated by the strong emission line.
@@ E.2 Emission lines in V2376 Ori b spectrum
We detect the [N ii] 6584, [O i] 6300, 6363 and [O ii] 7320, 7330 lines, which are commonly associated with jets and outflows (Solf & Boehm 1993; Hartigan et al. 1995; Flores-Rivera et al. 2023), as well as with externally irradiated disks (Haworth et al. 2023; Aru et al. 2024). Notably, we also detect the forbidden [C i] 8727 line–to our knowledge, the first such detection in a PMC. This line has recently been observed in irradiated disks in Orion (Haworth et al. 2023; Aru et al. 2024, 2024) and its excitation is attributed to far-UV (FUV) irradiation of neutral gas, either as a result of C+ recombinations (Escalante 1991) or FUV radiative pumping of neutral atomic carbon (Goicoechea et al. 2024). Detection of the Li absorption line is compatible with V2376 Ori b being a young, low-mass object (Magazzu et al. 1993; Phan-Bao et al. 2017). The presence of CaII (8662 ), HeI (6678, 7066 ), and OI (8447 ) lines points to a strongly accreting system (White & Basri 2003; Jayawardhana et al. 2003; Bowler et al. 2014). This is also in line with the strong emission in H, and HeI (Jayawardhana et al. 2003; White & Basri 2003; Eriksson et al. 2020). These accretion tracers have previously been detected in planetary-mass companions (Muzerolle et al. 2000; Natta et al. 2004; Joergens et al. 2013; Eriksson et al. 2020).
@@ Appendix G Accretion rate
The presence of several accretion tracers and the young age of the system ( Myr) strongly suggest that accretion is ongoing in V2376 Ori b. H recombination line fluxes can be used to estimate the associated accretion luminosities and rates. In Appendix G.1 we use the measured MUSE H luminosity of V2376 Ori b () combined with the results of the planet-surface shock model from Aoyama et al. (2021), and estimate an accretion luminosity of . In Appendix G.2 we also use the Pa luminosity derived from JWST data () combined with the empirical relation of Rogers et al. (2024) and find an accretion luminosity of . The accretion luminosities from both methods are in good agreement. We therefore consider a final accretion luminosity as the mean of those found with both methods, . Considering a radius and a mass (Table 1) this leads to a mass accretion rate .
G.1 From H emission observed with MUSE
We measure the flux by performing a Gaussian fit on the convolved spectrum using the tools provided in the MPDAF package developed by Piqueras et al. (2017). The measurement is based on the degraded-resolution spectrum shown in Fig. 2D, and yields a flux of . Considering the H luminosity derived in the degraded spectrum (Fig. 2 D) with a distance to the source of 304 pc, this leads to . To convert this luminosity into an accretion luminosity, we adopt the planetary surface shock model from Aoyama et al. (2021) that considers the contribution of shocks between the circumstellar disk and the circumplanetary disk to the total H emission:
This relation yields an accretion luminosity .
G.2 From Paschen F187N NIRCam filter
The accretion rate is related to the Paschen emission, traced by the F187N NIRCam filter. Chown et al. (2024) shows that the Paschen lines can be estimated with the equations:
with the band width and the flux density in the corresponding filter and Ap the aperture area in steradian used to measure the flux777in the F187N, the surface of the 2-pixels circular aperture is sr.. Considering the flux densities extracted in Sect. 3.1 at a distance of 304 pc and propagating the error on the distance and fluxes, we obtain a Paschen luminosity . Rogers et al. (2024) determined an empirical relation between the accretion luminosity and the Paschen luminosity for pre-main sequence stars:
Extrapolating this relation to the planetary-mass domain, this yields for V2376 Ori b an accretion luminosity .
@@ Appendix H Atmospheric model fitting to V2376 Ori b MUSE spectrum
Fig. 9 presents the best-fits models obtained for the V2376 Ori b MUSE spectrum using a blackbody emission, the ATMO2020 model (Phillips et al. 2020), and the BT-Cond model (Allard et al. 2001, 2011, 2012). Among these three models, the black-body (panel 9A) provides the best fit (and fewer number of free parameters) with a photosphere effective temperature of 1850 K. The ATMO2020 model (panel 9B) hardly reproduces the observed spectrum. Moreover, the best-fit yields an effective temperature of 2900 K for the source, which is difficult to reconcile with the observed luminosity and out of the range of the recommended temperatures for ATMO. The BT-Cond models (panel 9C), which are more suited for high-temperature, low-mass objects, better reproduce the spectrum than ATMO2020, but still present higher residual amplitudes than a simple blackbody. The effective temperature of 2900 K derived from this model appears significantly higher than the previous estimate made from the SED in Sect. 3.1 and implies an underestimated flux in the JWST filters as shown in Fig. 10. Overall, this comparison suggests that current atmospheric models are still not fully adapted to young, low-mass objects such as V2376 Ori b. This can be attributed to the limited number of such objects observed and thoroughly characterized to date.

### v2

@@ Appendix A Data presentation and reduction
We used the JWST NIRCam images of the ONC from the PDRs4All Early Release Science (ERS) program (Berné et al. 2022, PID 1288) that were previously presented in Habart et al. (2024). We selected the short-wavelength channel (i.e. for m), which includes seven filters (F140M, F162N, F164N, F182M, F187N, F210M, and F212N) with a 0.03” angular resolution. Details on the data reduction procedure to obtain these images are given in Habart et al. (2024). The images themselves are available through the specific PDRs4All MAST archive111https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html and through the REGARDS database222https://regards.osups.universite-paris-saclay.fr/user/jwst. We also use the Hubble Space Telescope (HST) images from the ACS/WFC of V2376 Ori presented in Ricci et al. (2008). This includes five filters (F435W, F555W, F658N, F775W, and F850LP), covering a wavelength range from 0.435 to 0.85 m.
The VLT MUSE observations were taken on February 27, 2025 (DDT P114 proposal 114.28JY). The observations originally consisted of four dithers, each with an integration time of 1800 sec and 90 degrees rotation between them. Due to a technical loss only the first dither position was successfully observed, yet, after inspection, the quality of this single exposure was deemed sufficient for this initial analysis.
The data were reduced using the ESO MUSE data reduction pipeline (v.2.10.14) together with the python wrapper, which is part of MUSEpack (Zeidler 2019; Zeidler et al. 2019). We apply all steps of the MUSE data reduction pipeline using the ESO provided master calibrations (MASTERBIAS, MASTERDARK, MASTERFLAT). To subtract the background we use the standard method ‘model’ using 10% of the datacube spaxels (fraction=10) after ignoring the darkest 10% (ignore=10). We visually inspect the skymask to ensure that it selects the darkest regions of our cube while rejecting all artifacts. Furthermore, we carefully inspect the sky spectrum and sky continuum to ensure that only telluric lines are removed from the final datacube. Once all data will be available we will combine these with the current dataset, which will further enhance our data product by allowing us to remove any remaining artifacts and significantly increase the signal-to-noise.
@@ B.1 Relative position between HST and JWST images
The proper motion of V2376 Ori provided by Gaia DR3 is ra=-0.148 mas/yr, dec=-1.082mas/yr. Even considering the largest temporal span of nearly 21 years between HST and VLT, the proper motion is expected to be smaller than one pixel, making common proper motion confirmation difficult. Therefore we only analyse the relative position of V2376 Ori and the companion at two different epochs using the HST F435W image (taken in October 2004) and JWST F187N (taken in September 2022) images. We exclude the MUSE image from this analysis because of the lack of additional sources in the field of view preventing robust astrometric alignment. Figure 4 shows the contours of the HST F435W (blue continuous lines) and JWST F187N (dashed red lines) images. Both are centered on V2376 Ori, and oversampled on a common grid to ensure alignment to about a tenth of a NIRCam pixel (0.003”). The companion (to the southwest of V2376 Ori) is shifted by = 9 mas, dec = 25 mas between the two epochs, i.e. over nearly 18 years. Considering that the pixel size is 50 mas for Hubble and 30 mas for NIRCam, and that the source shows a non-symmetric shape in the Hubble image, this difference is compatible with the absence of relative movement of the two sources over this time period.
@@ B.2 Angular separation of the sources
We estimated the angular separation between V2376 Ori and its companion in the HST F435W image, JWST F164N image and in the integrated MUSE cube. We first extracted an intensity profile along a cut performed along the axis connecting the two sources. We then determine the position of V2376 Ori and its companion in the profile with the highest and second highest intensity peak, respectively. The distance in pixels between these two peaks is finally converted into an angular separation considering the respective pixel sizes in arcsec of the different instruments333the pixel sizes for ACS/WFC HST, NIRCam JWST and VLT MUSE data are respectively 0.05”, 0.031”, and 0.025”. This yields for the F435W HST image, F164N JWST image and the MUSE image an angular separation of 0.230.05, 0.250.03” and 0.230.03”, respectively. The final angular separation is considered as the mean of these values, leading to = 0.230.03”.
@@ Appendix C Extraction of the spectral energy distribution
Figure 3 shows the SED of V2376 Ori b extracted from HST-ACS/WFC and JWST-NIRCam images. The flux arising from V2376 Ori b is contaminated by the background nebula emission from the ONC (see Fig. 1 C) and by the diffraction pattern (PSF) due to V2376 Ori (see, e.g. Fig. 1 F, G). The nebular emission is estimated using the Background2D function from the Photutils Python library444https://photutils.readthedocs.io/en/2.1.0/user_guide/index.htm. We use 5050 pixel boxes to create a low-resolution map with a sigma-clipping of three, which is then filtered with a 2D median filter of 55 pixels and interpolated to obtain the final 2D background. The images are then subtracted from this 2D background. Next we use three different approaches to remove the contribution of the diffraction spikes:
@@
For each image, we subtract a version rotated by 180° around the stellar center, so that every diffraction spike is cancelled by its counterpart on the opposite side of the star.
@@
Finally, the on-source flux in each background-subtracted and PSF corrected image is extracted using the AperturePhotometry function from the Photutils library with an aperture radius of 0.08” centered on V2376 Ori b, which is shown in Fig. 1 (G) with the dashed red circle. The final flux in a given filter is the mean of the fluxes obtained for the different PSF subtraction methods. Figure 5 presents the observed emissions at the position of the companion, at a position symmetric to that of the companion wrt to the star representing the contribution from the PSF (+background) at the companion position, and the emission from background nebula. The background nebula contributes to the on-source flux at a level of less than in most filters (except the F187N filter at 1.87 µm where this contribution is ). The stellar PSF contributes to at most of the on-source flux. This shows that the on-source flux is mainly affected by the stellar PSF rather than by the background nebular emission, but still largely dominated by emission from V2376 Ori b itself. Since the PSF and background contributions are both smaller than the flux of V2376 Ori b, we expect that residual emission after subtraction contributes to at most a few .
@@ Appendix D Physical properties of the V2376 Ori system
Table 1 summarizes the physical properties of the companion V2376 Ori b. Figure 6 shows the luminosity (top panel) and temperature (bottom panel) of V2376 Ori b, used to estimate its mass according to evolutionary tracks from Baraffe et al. (2003). The derivation of the companion’s luminosity is described in Sect. 3.1.
@@ E.1 Extraction of MUSE spectra
The spectrum of V2376 Ori and its companion are extracted from the MUSE cube. For the star, we use a circular aperture of 0.1” radius centered on the source. We then subtract the background nebula emission, estimated within two circular apertures of 0.1” radius, placed on either side of the star, in the direction perpendicular to the star-companion axis. For the companion, we perform the subtraction using two apertures situated on the sides of V2376 Ori b following an approach similar to Eriksson et al. (2020) for Delorme 1 (AB)b. The used apertures are shown on Fig. 8. The spectrum of V2376 Ori b is obtained by subtracting the on-source spectrum from the average background+stellar PSF contribution. Between 7950 and 8050 , the background nebula contributes to the on-source spectrum to %, whereas the PSF contributes to . Since the PSF and background contribution are both smaller than the flux of V2376 Ori b, we expect that residual emission after subtraction contributes to at most a few in the continuum. The mean of the two obtained background+PSF spectra is then subtracted from the on-source spectrum to obtain the final spectrum of the companion. The direct subtraction leads to artifacts with negative intensities and P-cygni-like profiles, possibly produced by velocity effects and/or thermal broadening. We apply a gaussian filter with a standard deviation of 2.5 pixels to remove these effects. The obtained spectra are then calibrated using the photometric fluxes derived in the same aperture in the F775W HST filter using the MPDAF Python package and associated filter throughput. This filter’s wavelength range is fully covered by MUSE and is not dominated by the strong emission line.
@@ E.2 Emission lines in V2376 Ori b spectrum
We detect the [N ii] 6584, [O i] 6300, 6363 and [O ii] 7320, 7330 lines, which are commonly associated with jets and outflows (Solf and Boehm 1993; Hartigan et al. 1995; Flores-Rivera et al. 2023), as well as with externally irradiated discs (Haworth et al. 2023; Aru et al. 2024b). Notably, we also detect the forbidden [C i] 8727 line–to our knowledge, the first such detection in a PMC. This line has recently been observed in irradiated discs in Orion (Haworth et al. 2023; Aru et al. 2024b, a) and its excitation is attributed to far-UV (FUV) irradiation of neutral gas, either as a result of C+ recombinations (Escalante 1991) or FUV radiative pumping of neutral atomic carbon (Goicoechea et al. 2024). Detection of the Li absorption line is compatible with V2376 Ori b being a young, low-mass object (Magazzu et al. 1993; Phan-Bao et al. 2017). The presence of CaII (8662 ), HeI (6678, 7066 ), and OI (8447 ) lines points to a strongly accreting system (White and Basri 2003; Jayawardhana et al. 2003; Bowler et al. 2014). This is also in line with the strong emission in H, and HeI (Jayawardhana et al. 2003; White and Basri 2003; Eriksson et al. 2020). These accretion tracers have previously been detected in PMCs (Muzerolle et al. 2000; Natta et al. 2004; Joergens et al. 2013; Eriksson et al. 2020).
@@ Appendix G Accretion rate
The presence of several accretion tracers and the young age of the system ( Myr) strongly suggest that accretion is ongoing in V2376 Ori b. We measure the flux by performing a Gaussian fit on the convolved spectrum using the tools provided in the MPDAF package developed by Piqueras et al. (2017). The measurement is based on the degraded-resolution spectrum shown in Fig. 2D, and yields a flux of . Considering this H flux with a distance to the source of 304 pc, this leads to . To convert this luminosity into an accretion luminosity, we adopt the planetary surface shock model from Aoyama et al. (2021) that considers the contribution of shocks between the circumstellar disc and the circumplanetary disc to the total H emission:
This relation yields an accretion luminosity . Considering a radius and a mass (Table 1) this leads to a mass accretion rate .
@@ Appendix H Atmospheric model fitting to V2376 Ori b MUSE spectrum
Figure 11 presents the best-fits models obtained for the V2376 Ori b MUSE spectrum using a blackbody emission, the ATMO2020 model (Phillips et al. 2020), and the BT-Cond model (Allard et al. 2001, 2011, 2012). Among these three models, the black-body (panel 11A) provides the best fit (and fewer number of free parameters) with a photosphere effective temperature of 1850 K. The ATMO2020 model (panel 11B) hardly reproduces the observed spectrum. Moreover, the best-fit yields an effective temperature of 2900 K for the source, which is difficult to reconcile with the observed luminosity and out of the range of the recommended temperatures for ATMO. The BT-Cond models (panel 11C), which are more suited for high-temperature, low-mass objects, better reproduce the spectrum than ATMO2020, but still present higher residual amplitudes than a simple blackbody. The effective temperature of 2900 K derived from this model appears significantly higher than the previous estimate made from the SED in Sect. 3.1 and implies an underestimated flux in the JWST filters as shown in Fig. 12. Overall, this comparison suggests that current atmospheric models are still not fully adapted to young, low-mass objects such as V2376 Ori b. This can be attributed to the limited number of such objects observed and thoroughly characterized to date.